value of cos(22.5degree) in surd form..???

Grade:11

3 Answers

APURV GOEL
12 years ago

we have 1+cos2x = 2(cosx)2

cosx = root(1/2 + cos2x/2)

cos22.5 = root(1/2 + cos45/2)

cos22.5 = (1/2+cos45/2)1/2

Abhishek kumar
5 years ago
Cos(2×22.5)=2cos^2A-1cos45=2cos^2(22.5)-11/√2=2cos^2(22.5)-1[cos45=1/√2]1/√2+1=cos^2(22.5)√2+1/2√2=cos^2(22.5)[rationalise do]√2+2/2=cos22.5
Soumendu Majumdar
5 years ago
Dear Student,
1+cos 45^{\circ}= 2 cos^222.5^{\circ}    using 1 + cos 2A = 2 cos^2A
so    cos^222.5^{\circ}= (\sqrt{2}+ 1)/2\sqrt{2}
so   cos22.5^{\circ}= \sqrt{(\sqrt{2}+ 1)/2\sqrt{2}}     
 
Hope it helps!
with regards,
Soumendu
P.S- If you have any doubt feel free to ask, I’ll be glad to help you!